# Prompting Large Language Models for Recommender Systems: A Comprehensive Framework and Empirical Analysis

### v1

@@ 4.1.3. Model Architecture
For the architecture design of LLMs leveraged as recommender systems, we consider three mainstream architectures as summarized in (Zhao et al., 2023), i.e., encoder-decoder, prefix decoder, and causal decoder. The recommendation scenarios for each framework are introduced in what follows.
$\bullet$ The encoder-decoder architecture adopts two Transformer blocks as the encoder and decoder, which is the basis of BERT (Vaswani et al., 2017). In line with existing literature on LLM-based recommender systems, the bidirectional property of the encoder-decoder architecture allows LLMs to easily customize encoders and decoders towards recommendation (e.g., dual-encoder considering both ids and text (Qiu et al., 2023)), and conveniently adapt to multiple recommendation tasks (Chu et al., 2023; Li et al., 2023j; Geng et al., 2022; Zhang et al., 2023f; Qiu et al., 2023). A few LLMs use the encoder-decoder architecture, and the typical one is the series of T5 and its variant Flan-T5 (Chung et al., 2022).
$\bullet$ The prefix decoder architecture is also the decoder-only architectures, which is known as non-causal decoder. It can bidirectionally encode the prefix tokens like the encoder-decoder architecture, and perform unidirectional attention on the generated tokens like the causal decoder architecture. One of the representative LLMs based on the prefix decoder architecture is ChatGLM (Zeng et al., 2022) and its variants, and researchers have attempted to explore the recommendation performance of ChatGLM as one of the benchmarks in related work (Liu et al., 2023d).
$\bullet$ The causal decoder architecture has been widely adopted in various LLMs, and the series of GPT (Brown et al., 2020) and LLaMA (Touvron et al., 2023a) are the most representative models. It uses the unidirectional attention mask, and only decoders are deployed to process both the input and output tokens. Due to the popularity of the causal decoder architecture, most LLM-based recommender systems employ this framework to adapt to different recommendation tasks such as click-through rate predictions, sequential recommendation and conversational recommender systems.
@@ 4.1.4. Parameter Scale
To meet the diverse needs of different users, the most typical variant of LLMs is the parameter scale. In general, open-source models have multiple parameter sizes to choose from, and larger parameter sizes generally mean better capabilities (Touvron et al., 2023a; Zhao et al., 2023). But meanwhile, the corresponding computational and spatial complexity will also increase. Considering the memory and efficiency issues for experiments, researchers in the field of LLM-based recommender systems generally use LLMs with parameters no more than 10B, while the performance of LLMs with larger parameters remains to be further explored in the field of recommender systems (Kang et al., 2023).
@@ 4.1.5. Context Length
Another property closely related to user needs is the length of the input context. The inability to handle inputs with longer contexts means that decision-making cannot be accurately developed, thereby limiting the model capabilities (Zhao et al., 2023; Touvron et al., 2023b). When the user input exceeds the limited length, the input will be truncated, so sufficient context length is crucial for the user experience (Hou et al., 2023; Dai et al., 2023a; Ma et al., 2023). However, different lengths of context inputs imply different model architectures and parameters. When expanding the context length of a model, it often leads to higher time and memory complexity. To address the length limitation of LLMs, existing methods either selectively discard previous contexts using sliding windows (Xiao et al., 2023), or only sample a portion of the context for retrieval augmentation (Lewis et al., 2020; Lin et al., 2023b), or employ small models without emergence ability. Despite recent strategies, the limitation of context length has not yet been truly resolved. Considering economic and efficiency issues, existing LLMs including open-source and closed-source models only provide a limited number of context length options. In this paper, we mainly focus on several classic lengths, i.e., 2K, 4K, 8K, 16K and 32K (K is the abbreviation for one thousand, similarly hereinafter).
@@ 4.2.1. Research Questions
In this section, we conduct experiments to verify the impact of different factors of LLMs on recommendation results. Specifically, we focus on the following four research questions.
  • RQ1: What are the differences between the recommendation ability of LLMs and traditional recommenders?
  • RQ2: How do the different attributes of LLMs, including the public availability, model architectures, parameter scales, and context lengths affect the recommendation performance and inference time?
  • RQ3: What are the similarities and differences in recommendation results of LLMs with different tuning strategies? Is the LLM after instruction tuning more suitable for recommendation tasks?
@@ 4.2.2. Evaluated Models
As for the experimental settings, we consider the following baselines and LLMs.
@@
  • Pop: Pop method always ranks the candidate items based on their interaction times with users in the training set. We consider it as the fully-trained method since it uses the statistical information of datasets.
@@
  • SASRec (Kang and McAuley, 2018): SASRec is a sequential recommendation model based on the backbone of the classic self-attention network a.k.a., Transformer (Vaswani et al., 2017), and achieves comparable performance among sequential models.
@@
  • ChatGPT: ChatGPT is a closed-source large-scale pre-trained language model developed by OpenAI. Note that OpenAI has released interfaces of ChatGPT on March 1 and June 13, 2023, respectively. Considering the up-to-date requirements, we adopt the version on June 13 for recency, the same as GPT-4.
  • GPT-4: GPT-4 is the latest generation of closed-source natural language processing model launched by the OpenAI company. Experiments have shown that GPT-4 is significantly superior to ChatGPT in multiple tasks.
@@
  • Flan-T5 (Chung et al., 2022): Flan-T5 is an open-source language model based on the encoder-decoder architecture T5 released by Google (Raffel et al., 2020). Flan-T5 is extended from T5 by a multi-task fine-tuning paradigm i.e., instruction tuning to enhance the generalization of different tasks. There are multiple variants of Flan-T5 in terms of parameters, including Flan-T5-Small (80M), Flan-T5-Base (250M), Flan-T5-Large (780M), Flan-T5-XL (3B) and Flan-T5-XXL (11B). Since the first three models are too small to meet the requirements of LLMs discussed in this paper (1B, B is short for billion and the same below), we consider the Flan-T5-XL and Flan-T5-XXL for comparison.
@@
  • ChatGLM (Zeng et al., 2022): ChatGLM is an open-source bilingual dialogue LLM that supports both Chinese and English, based on the General Language Model (GLM) (Zeng et al., 2022) with the prefix decoder architecture. The team released the second version ChatGLM2 and the third version ChatGLM3 in June 2023 and October 2023, respectively.
@@
  • LLaMA (Touvron et al., 2023a): LLaMA is an open-source language model introduced by MetaAI from the causal decoder architecture with four sizes (6B, 13B, 35B and 65B). Due to its outstanding performance and low computational cost, LLaMA has received much attention from researchers so far, and Vicuna (Chiang et al., 2023) is one of the most popular variants by extending LLaMA. To further improve abilities of LLaMA, MetaAI released LLaMA2 as the next generation of open-source large language models in July, 2023. In addition to the regular version, MetaAI also provide the chatting version of LLaMA2 (i.e., LLaMA2-chat) (Touvron et al., 2023b), and it is specifically tuned for the dialogue scenario by Reinforcement Learning with Human Feedback (RLHF).
Table 4. Overall performance of different models on recommendation. We consider both the fully-trained settings for traditional models and zero-shot settings for LLMs. Note that there are always ground-truth items in the randomly selected 20 candidates, so the ideal recall@20 equals to 1. “IT” is for “Inference Time”, and we record the average inference time for each user measured in seconds (s). “N/A” is the abbreviation for “Not Applicable” since the inference time of closed-source models is unknown.
MovieLens-1M Amazon-Books
model context length param. size recall@20 ndcg@1 ndcg@10 IT (s) recall@20 ndcg@1 ndcg@10 IT (s)
@@
Open-source LLMs with the causal decoder architecture
@@
LLaMA2 (chat) 4K 70B 0.9550 0.2430 0.4344 23.07 0.8850 0.2230 0.3827 25.01
@@ 4.3.1. LLMs Compared to Traditional Recommenders (RQ1)
$\bullet$ Recommendation performance of LLMs. As shown in Table 4, we provide the fully-trained results of four traditional methods, as well as the zero-shot recommendation performance of various LLMs. For traditional recommenders, BPR (Rendle et al., 2009) based on collaborative filtering is significantly better than Pop based on popularity, and the sequential recommendation model SASRec (Kang and McAuley, 2018) combined with attention mechanism and temporal information is significantly better than BPR, which is consistent with the results in existing literature (Rendle et al., 2009; Kang and McAuley, 2018; Zhou et al., 2020). It is worth noting that for the 20 candidate items, LLMs that rely on natural languages cannot completely recall all items in most cases, and several models with poor abilities can only output a dozen items, greatly limiting the accuracy of recommendation results. Therefore, recall@20 indicates the ability of LLMs to memorize, re-rank and output candidate items, and several approaches such as the re-generation method (Li et al., 2023i) and probability distribution outputs (Yue et al., 2023) have been proposed to improve the recall performance. While for the traditional models, all candidate items can be recalled (i.e., recall@20 equals to 1). Because the candidate items are not recalled completely, the recommendation effect of a few LLMs (e.g., Flan-T5 and ChatGLM) is not even as good as the random baseline. For most LLMs, the zero-shot recommendation performance is not as good as the baseline method Pop based on popularity of interactions in the dataset (Liu et al., 2023c; Hou et al., 2023; Dai et al., 2023a). However, the powerful LLMs like ChatGPT and LLaMA-70B (chat) (Touvron et al., 2023b) can achieve better results than Pop in zero-shot settings. Furthermore, GPT-4 can even perform better than the fully-trained matrix factorization model BPR on two datasets, indicating the potential of LLMs to serve as the backbone of recommender systems. In addition, the significant differences between the results of LLMs demonstrate the importance of selecting appropriate LLMs for downstream recommendation tasks (Kang et al., 2023; Liu et al., 2023d, c).
@@
(b) Amazon-Books
Figure 3. The recommendation performance of LLMs w.r.t. the number of historical item sequences.
$\bullet$ The impact of historical item sequences. As for ranking tasks in Fig. 2(a), the recently interacted historical items are used as the user representations. However, there is no standard value for the number of items that represent users. To analyze the impact of historical item sequences on the recommendation performance, we conduct experiments to explore the recommendation effect of the sequential recommender SASRec (Kang and McAuley, 2018) and the closed-source LLM ChatGPT with different numbers of historical interactions. For the fully-trained SASRec, the maximum length of the historical item sequence will affect the model framework and prediction results (Kang and McAuley, 2018; Hou et al., 2022). In order to ensure the model uniformity, we fix the model checkpoint with the historical item sequence of 50 as the “SASRec (fixed)” for comparison, and evaluate the recommendation performance (NDCG@10 (Järvelin and Kekäläinen, 2002)) with the number of historical items at 1, 5, 10, 20, 30, 40 and 50, respectively. For ChatGPT and GPT-4, we get zero-shot results with different items to verify whether the powerful LLMs can deal with the long context for recommendation. As illustrated in Fig. 3, with the increasing number of historical items, the results of SASRec improve steadily, while the the recommendation performance of LLMs changes little. Consistent conclusions on two datasets can be drawn that even if LLMs can accept more historical items for user representations, increasing the number of historical items does not bring significant gains in the recommendation performance. The performance trends of LLMs show that the increased historical item sequence is not fully utilized by the language model, indicating the importance of selecting appropriate item sequences to represent users, and the inadequacy of LLMs for user interest mining. To improve the mining of user interest for LLMs, approaches such as retrieval augmentation (Lin et al., 2023b) and prompting strategies (Wang et al., 2023c; Yao et al., 2023) can be used, which will be analyzed in Section 5.
@@
Observations on the Overall Recommendation Performance of LLMs In zero-shot scenarios, LLMs have cold-start capabilities, and GPT-4 even surpasses collaborative filtering models. However, all LLMs are inferior to fully-trained sequential recommendation models. Even if LLMs can accept more historical items for user representations, increasing the number of historical items does not bring significant gains in the recommendation performance. Compared to recommenders, the inference time of LLMs is unacceptable to be used for real applications.
@@
$\bullet$ Public availability. As shown in Table 4, closed-source models achieve significantly better results than the open-source models in the cold-start scenario, but they cannot outperform fully-trained sequential models. In terms of LLMs, ChatGPT with zero-shot settings has comparable recommendation performance with the fully-trained Pop especially on the sparse Amazon-Books dataset, indicating the fundamental ability of LLMs on recommendation tasks. Furthermore, the upgraded GPT-4 exceeds ChatGPT by a large margin due to its strong zero-shot generalization ability. The superior zero-shot performance of GPT sheds lights on leveraging LLMs for recommendation. However, the open-source models always get poor results compared to GPT-4 in the zero-shot settings, while LLaMA2-chat-70B has the comparable recommendation performance with ChatGPT. The reason is that open-source models lack comprehensive cold-start capabilities, and their strength lies in the ability to integrate domain knowledge through strategies such as prompt tuning. In line with previous studies on LLMs (Hou et al., 2023; Kang et al., 2023; Ma et al., 2023; Liu et al., 2023c), employing a closed-source model in cold-start scenarios yields better results, while an open-source model is more flexible and easy to use when tuning is needed (Bao et al., 2023b; Zheng et al., 2023; Li et al., 2023b; Luo et al., 2023b).
@@
$\bullet$ Model architecture. For the model architecture, Flan-T5 (Chung et al., 2022) based on the encoder-decoder architecture has almost no ability to recommend items in the cold-start setting, as its training corpus does not involve specialized instructions of our task. Trained with prompts of recommendations, the encoder-decoder architecture is suitable for prompt tuning and instruction tuning (Zhang et al., 2023f; Geng et al., 2022; Chu et al., 2023). Similarly, the first and second versions of ChatGLM (Zeng et al., 2022) based on the prefix decoder perform poorly on the zero-shot ranking task, and are not as good as Vicuna and LLaMA2 based on the causal decoder. However, the third version of ChatGLM, i.e., ChatGLM3 has comparable recommendation performance with Vicuna (Chiang et al., 2023) and LLaMA2 (Touvron et al., 2023b), which further indicates the importance of selecting an advanced foundation model. In terms of the series of LLaMA models (Touvron et al., 2023a), LLaMA2 is better than Vicuna, and Vicuna is better than LLaMA, which is related to their training data and release time. Furthermore, the chat version of LLaMA2, i.e., LLaMA2-chat is a series that uses conversational dialog instructions to fine-tune LLaMA2 (Touvron et al., 2023b), which is more suitable for our tasks in ranking settings. Therefore, the results of LLaMA2-chat are significantly improved compared with LLaMA2, and LLaMA2-70B-chat even achieves better performance than the closed-source model ChatGPT. Generally speaking, researchers prefer to study recommendation tasks based on the causal decoder framework such as LLaMA, and the second version of LLaMA has a better generalization ability than the first version in recommendation tasks.
@@
(b) Amazon-Books
Figure 4. The recommendation performance and inference time of LLMs w.r.t. the parameter scale. A larger point means more parameters.

### v2

@@ 4.1.3. Model Architecture
For the architecture design of LLMs leveraged as recommender systems, we consider three mainstream architectures as summarized in (Zhao et al., 2023), i.e., encoder-decoder, prefix decoder, and causal decoder. The recommendation scenarios for each framework are introduced.
$\bullet$ The encoder-decoder architecture adopts two Transformer blocks as the encoder and decoder, which is the basis of BERT (Devlin et al., 2019). In line with existing literature on LLM-based recommender systems, the bidirectional property of the encoder-decoder architecture allows LLMs to easily customize encoders and decoders towards recommendation (e.g., dual-encoder considering both ids and text (Qiu et al., 2023)), and conveniently adapt to multiple recommendation tasks (Chu et al., 2023; Li et al., 2023i; Geng et al., 2022; Zhang et al., 2023e; Qiu et al., 2023).
$\bullet$ The prefix decoder architecture is also the decoder-only architectures, which is known as non-causal decoder. It can bidirectionally encode the prefix tokens like the encoder-decoder architecture, and perform unidirectional attention on the generated tokens like the causal decoder architecture. One of the representative LLMs based on the prefix decoder architecture is ChatGLM (AI, 2024), and researchers have attempted to explore the recommendation performance of ChatGLM as one of the benchmarks in related work (Liu et al., 2023d). Note that the variants of ChatGLM, i.e., ChatGLM2 and ChatGLM3 employ the causal decoder architecture.
$\bullet$ The causal decoder architecture has been widely adopted in various LLMs, and the series of GPT (Brown et al., 2020) and LLaMA (Touvron et al., 2023a) are the most representative models. It uses the unidirectional attention mask, and only decoders are deployed to process both the input and output tokens. Due to the popularity of the causal decoder architecture, most LLM-based recommender systems employ this framework to adapt to different recommendation tasks such as CTR predictions and sequential recommendation.
@@ 4.1.4. Parameter Scale
To meet the diverse needs of different users, the most typical variant of LLMs is the parameter scale. In general, open-source models have multiple parameter sizes to choose from, and larger parameter sizes generally mean better capabilities (Touvron et al., 2023a; Zhao et al., 2023). But meanwhile, the corresponding computational and spatial complexity will also increase. Considering the memory and efficiency issues for experiments, researchers in the field of LLM-based recommender systems generally use LLMs with parameters no more than 10B (B is short for billion and the same below), while the performance of LLMs with larger parameters remains to be further explored in the field of recommender systems (Kang et al., 2023).
@@ 4.1.5. Context Length
Another property closely related to user needs is the length of the input context. The inability to handle inputs with longer contexts means that decision-making cannot be accurately developed, thereby limiting the model capabilities (Zhao et al., 2023; Touvron et al., 2023b). When the input exceeds the limited length, the input will be truncated, so sufficient context length is crucial for the user experience (Hou et al., 2024; Dai et al., 2023a; Ma et al., 2023). However, different lengths of context inputs imply different model architectures and parameters. When expanding the context length of a model, it often leads to higher time and memory complexity. To address the length limitation of LLMs, existing methods either selectively discard previous contexts using sliding windows (Xiao et al., 2024), or only sample a portion of the context for retrieval augmentation (Lewis et al., 2020; Lin et al., 2024a), or employ small models without emergence ability. Despite recent strategies, the limitation of context length has not yet been truly resolved. Considering economic and efficiency issues, existing LLMs including open-source and closed-source models only provide a limited number of context length options. In this paper, we mainly focus on several classic lengths, i.e., 2K, 4K, 8K, 16K, 32K and 128K (K is the abbreviation for one thousand, the same below).
@@ 4.2.1. Research Questions
In this section, we conduct experiments to verify the impact of different factors of LLMs on recommendation results focusing on the following four Research Questions (RQ).
  • RQ1: What are the performance differences between LLMs and traditional recommenders?
  • RQ2: How do different attributes of LLMs, including the public availability, model architectures, parameter scales, and context lengths affect the performance and inference time?
  • RQ3: What are the similarities and differences in recommendation results with different tuning strategies? Is the LLM after instruction tuning more suitable for recommendation?
@@
4.2.2. Experimental Setup
This section considers two prevalent recommendation scenarios based on the tuning status of LLMs. To address RQ1, RQ2, and RQ4, we evaluate the zero-shot ranking efficacy of LLMs by using prompts as illustrated in Fig. 2(a). In the context of RQ3, we fine-tune LLaMA-based LLMs to act as CTR predictors, with the corresponding data sample presented in Fig. 2(b). Details regarding the experimental configurations are available in Section 3.3.
@@ 4.2.3. Evaluated Models
For empirical analysis, we assess the following baselines and LLMs.
@@
  • Pop: Pop ranks the candidate items based on their interaction times in the training set. We consider it as the fully-trained method since it uses the statistical information of datasets.
@@
  • SASRec (Kang and McAuley, 2018): SASRec is a sequential recommendation model based on the backbone of the classic self-attention network a.k.a., Transformer (Vaswani et al., 2017), and achieves comparable performance among sequential models. It utilizes the unidirectional attention mask for model training and achieves comparable performance among sequential models.
@@
  • ChatGPT: ChatGPT is a closed-source large-scale pre-trained language model developed by OpenAI, and we adopt the version on June 13, 2023, the same as GPT-4. We explore two context lengths of ChatGPT, i.e., 4K and 16K.
  • GPT-4 (OpenAI, 2023): GPT-4 is a closed-source LLM with the context length of 8K. Building upon the success of ChatGPT, represents the next evolutionary leap in LLMs by OpenAI, delivering unprecedented capabilities in natural language processing.
  • GPT-4o (OepnAI, 2024): GPT-4o is a state-of-the-art closed-source LLM developed by OpenAI, released in 2024. It features a 128K token context window, significantly enhancing its ability to handle extended conversations and complex tasks requiring long-term memory.
  • GPT-4o mini: GPT-4o mini is a cost-efficient version of the GPT-4o model (OepnAI, 2024), designed to deliver high-performance natural language processing capabilities with reduced computational requirements.
  • OpenAI o1 mini: OpenAI o1 mini is a closed-source lightweight reasoning model designed for applications that prioritize reasoning over broad factual knowledge, offering faster response time than GPT-4o. Both GPT-4o mini and OpenAI o1 mini support a 128K token context window.
@@
  • Flan-T5 (Chung et al., 2024): Flan-T5 is an open-source language model based on the encoder-decoder architecture T5 released by Google (Raffel et al., 2020). Flan-T5 is extended from T5 by a multi-task fine-tuning paradigm i.e., instruction tuning to enhance the generalization of different tasks. There are multiple variants of Flan-T5 in terms of parameters, including Flan-T5-Small (80M), Flan-T5-Base (250M), Flan-T5-Large (780M), Flan-T5-XL (3B) and Flan-T5-XXL (11B). Since the first three models are too small to meet the requirements of LLMs discussed in this paper (1B, B is short for billion and the same below), we consider the Flan-T5-XL and Flan-T5-XXL for comparison. All variants of Flan-T5 have a 512-token context length limit.
@@
  • ChatGLM (Zeng et al., 2023): ChatGLM is an open-source bilingual dialogue LLM that supports both Chinese and English, based on the General Language Model (GLM) (Zeng et al., 2023) with the prefix decoder architecture. The team released the second version ChatGLM2 and the third version ChatGLM3 in June 2023 and October 2023, respectively. Note that ChatGLM2 and ChatGLM3 have replaced the model architecture with a casual encoder. All three versions of open-source ChatGLM have a parameter scale of 6B. As for the context length, ChatGLM-6B supports 2K tokens while ChatGLM2-6B can cover 32K tokens at one time. Moreover, ChatGLM3-6B has both 2K and 32K versions.
@@
  • LLaMA (Touvron et al., 2023a): LLaMA is an open-source language model introduced by MetaAI from the causal decoder architecture with four sizes (7B, 13B, 33B and 65B). Due to its outstanding performance and low computational cost, LLaMA has received much attention from researchers so far.
  • Vicuna (Chiang et al., 2023): Vicuna (Chiang et al., 2023) is one of the most popular variants by extending LLaMA , and it has a context length of 4K tokens with two parameter scales (7B and 13B).
  • LLaMA2 (Touvron et al., 2023b): LLaMA2 is the next generation of LLaMA released in July, 2023. We also conduct experiments of the chatting version of LLaMA2 (i.e., LLaMA2-chat) (Touvron et al., 2023b), and it is specifically tuned for the dialogue scenario by Reinforcement Learning with Human Feedback (RLHF). Both LLaMA2 and LLaMA2-chat have three parameter scales (7B, 13B and 70B), and the context length is upgraded from 2K in LLaMA to 4K in LLaMA2.
  • LLaMA3 (Dubey et al., 2024): LLaMA3 is an evolution in the LLaMA series developed by Meta, released in 2024. It features a dense Transformer (Vaswani et al., 2017) architecture with up to 405 billion parameters and supports a context window of 128K tokens, enabling it to handle long and complex text sequences effectively. We mainly use two parameter scales of LLaMA3 (8B and 70B).
  • Qwen2.5 (Team, 2024): Qwen2.5 is a highly advanced series of LLMs developed by Alibaba, featuring a context length of up to 128K tokens and supporting outputs of up to 8K tokens, with model sizes ranging from 0.5B to 72B parameters, and we employ three parameter scales (7B, 32B and 72B).
  • DeepSeek-V3 (DeepSeek-AI, 2024): DeepSeek-V3 is an open-source LLM developed by DeepSeek, featuring a Mixture-of-Experts (MoE) architecture with 671 billion total parameters, of which 37 billion are activated per token. It also supports a 128K token context window.
  • DeepSeek-R1 (DeepSeek-AI, 2024): DeepSeek-R1 is a reasoning model developed by DeepSeek with long-chain reasoning capabilities and transparent slow-thinking process. DeepSeek-R1 is expected to be open-sourced in subsequent iterations, and we treat it as an open-source model in Table 4. The results are derived from calls made through the official platform  [^3] The URL of DeepSeek API: https://chat.deepseek.com/.
[…]
MovieLens-1M Amazon Books
Model Context Length Param. Size Coverage NDCG@1 NDCG@10 IT (s) Coverage NDCG@1 NDCG@10 IT (s)
@@
Open-source LLMs with the Causal Decoder Architecture
@@
LLaMA2 (chat) 4K 70B 0.9550 0.2430 0.4344 23.07 0.7850 0.2100 0.3827 25.01
8B 0.9250 0.1850 0.4234 13.98 0.9600 0.1750 0.3331 14.04
LLaMA3 128K 70B 1.0000 0.3100 0.5385 22.29 0.9900 0.2300 0.4144 24.57
7B 0.6300 0.0450 0.2793 2.72 0.6250 0.0400 0.2654 4.65
[…]
DeepSeek-R1 128K 671B 0.9850 0.3000 0.5270 17.08 0.9800 0.2350 0.4321 24.33
@@ 4.3.1. LLMs Compared to Traditional Recommenders (RQ1)
$\bullet$ Recommendation performance of LLMs. As shown in Table 4, we provide the fully-trained results of four traditional methods, as well as the zero-shot recommendation performance of various LLMs. For traditional recommenders, BPR (Rendle et al., 2009) based on collaborative filtering is significantly better than Pop based on popularity, and the sequential recommendation model SASRec (Kang and McAuley, 2018) combined with attention mechanism and temporal information is significantly better than BPR, which is consistent with the results in existing literature (Rendle et al., 2009; Kang and McAuley, 2018; Zhou et al., 2020). It is worth noting that for the 20 candidate items, LLMs that rely on natural languages cannot completely recall all items in most cases, and several models with poor abilities can only output a dozen items, greatly limiting the accuracy of recommendation results. Therefore, “coverage (recall@20)” indicates the ability of LLMs to memorize, re-rank and output candidate items, and several approaches such as the re-generation method (Li et al., 2023h) and probability distribution outputs (Yue et al., 2023) have been proposed to improve the recall performance. While for the traditional models, all candidate items can be recalled (i.e., coverage equals to 1). Because the candidate items are not recalled completely, the recommendation effect of a few LLMs (e.g., Flan-T5 and ChatGLM) is not even as good as the random baseline. DeepSeek-R1 performs worse than DeepSeek-V3, indicating that the long-chain reasoning capabilities and slow-thinking process of LLMs are optimized for Mathematics and coding tasks, which is not necessarily suitable for the zero-shot recommendation scenario. It might be perplexing to observe that the Qwen2.5 model (Team, 2024) performs well on the NDCG@10 metric, yet all three parameter scales yield equally poor results on NDCG@1. This discrepancy arises because Qwen2.5 tends to provide lengthy explanations in its output, which significantly reduces the top-1 accuracy during output parsing. When evaluating different versions of the same model, a general trend is observed where newer versions tend to outperform their predecessors. For instance, LLaMA3 (Dubey et al., 2024) typically surpasses LLaMA2 (Touvron et al., 2023b), and ChatGLM3 (AI, 2024) is generally superior to ChatGLM2. However, anomalies have been noted that ChatGLM outperforms ChatGLM2 on the Amazon Books dataset, and LLaMA-7B achieves a higher NDCG@10 score compared to LLaMA2-7B on the MovieLens-1M dataset. The underlying reason is that LLMs with parameters less than 10 billion often struggle to comprehend the prompts in recommendation tasks adequately, frequently resulting in outputs that do not meet the required format. We will explore this issue in detail through case studies in Section 4.3.4. For most LLMs, the zero-shot recommendation performance is not as good as the baseline method Pop based on popularity of interactions in the dataset (Liu et al., 2023c; Hou et al., 2024; Dai et al., 2023a). However, the powerful LLMs like ChatGPT and LLaMA-70B (chat) (Touvron et al., 2023b) can achieve better results than Pop in zero-shot settings. Furthermore, GPT-4 and GPT-4o can perform better than the fully-trained matrix factorization model BPR on the MovieLens-1M dataset, and DeepSeek-V3 outperforms BPR on two datasets, indicating the potential of LLMs to serve as the backbone of recommender systems. In addition, the significant differences between the results of LLMs demonstrate the importance of selecting appropriate LLMs for downstream recommendation tasks (Kang et al., 2023; Liu et al., 2023d, c).
@@
(b) Amazon Books
Figure 3. The recommendation performance of LLMs w.r.t. the number of historical item sequences. As in 2(a), we use recently interacted items to model user interest in the zero-shot ranking task. To compare the impact of historical item sequences on LLMs, we illustrate the ranking performance of the traditional sequential recommender SASRec (Kang and McAuley, 2018) and three closed-source LLMs with different numbers of historical items.
$\bullet$ The impact of historical item sequences. As for ranking tasks in Fig. 2(a), the recently interacted historical items are used as the user representations. However, there is no standard value for the number of items that represent users. To analyze the impact of historical item sequences on the recommendation performance, we conduct experiments to explore the recommendation effect of the sequential recommender SASRec (Kang and McAuley, 2018) and the closed-source LLM ChatGPT with different numbers of historical interactions. For the fully-trained SASRec, the maximum length of the historical item sequence will affect the model framework and prediction results (Kang and McAuley, 2018; Hou et al., 2022). To ensure the model uniformity, we fix the model checkpoint with the historical item sequence of 50 as the “SASRec (fixed)” for comparison, and evaluate the recommendation performance (NDCG@10 (Järvelin and Kekäläinen, 2002)) with different numbers of historical items. For ChatGPT and GPT-4, we get zero-shot results with different items to verify whether the powerful LLMs can deal with the long context for recommendation. In Fig. 3, with the increasing number of historical items, the results of SASRec improve steadily, while the performance of LLMs changes little. Consistent conclusions on two datasets can be drawn that even if LLMs can accept more historical items for user representations, increasing the number of historical items does not bring significant gains in the recommendation performance. The performance trends of LLMs show that the increased historical item sequence is not fully utilized by the language model. This phenomenon may be attributed to the fact that the long-sequence capabilities of LLMs have not been specifically optimized for zero-shot recommendation tasks. If targeted optimizations were applied to enhance the ability of LLMs to handle task-specific prompts, it is reasonable to expect that improvements in performance could be achieved. To further improve the mining of user interest for LLMs, approaches such as retrieval augmentation (Lin et al., 2024a) and prompting strategies (Wang et al., 2024c; Yao et al., 2023) can be used, which will be analyzed in Section 5.
@@
Observations on the Overall Recommendation Performance of LLMs In zero-shot scenarios, LLMs have cold-start capabilities, and GPT-4o even surpasses collaborative filtering models. However, all LLMs are inferior to fully-trained sequential recommendation models. The long-chain reasoning capabilities and slow-thinking process of LLMs are optimized for Mathematics and coding tasks, which is not necessarily suitable for the zero-shot recommendation scenario. Even if LLMs can accept more historical items for user representations, increasing the number of historical items does not bring significant gains in the recommendation performance, while traditional recommenders can better utilize longer sequences. It also shows that the user interest should be effectively modeled rather than blindly stacking historical items. The inference time of LLMs for recommendation is unacceptable for applications.
@@
$\bullet$ Public availability. As shown in Table 4, closed-source models achieve significantly better results than the open-source models in the cold-start scenario, but they cannot outperform fully-trained sequential models. The results of the closed-source LLMs can be further improved through approaches like retrieval augmentation, which we will discuss in Section 5.2. In terms of LLMs, ChatGPT with zero-shot settings has comparable recommendation performance with the fully-trained Pop especially on the sparse Amazon Books dataset, indicating the fundamental ability of LLMs on recommendation tasks. Furthermore, the upgraded GPT-4 and GPT-4o exceed ChatGPT by a large margin due to its strong zero-shot generalization ability. The superior zero-shot performance of GPT sheds lights on leveraging LLMs for recommendation. Although DeepSeek-V3 has the comparable recommendation performance with GPT-4o, the open-source models usually get poor results compared to GPT-4o in the zero-shot settings. The reason is that open-source models lack comprehensive cold-start capabilities, and their strength lies in the ability to integrate domain knowledge through strategies such as prompt tuning. In line with previous studies on LLMs (Hou et al., 2024; Kang et al., 2023; Ma et al., 2023; Liu et al., 2023c), employing a closed-source model in cold-start scenarios yields better results, while an open-source model is more flexible and easy to use when tuning is needed (Bao et al., 2023b; Zheng et al., 2024; Li et al., 2023c; Luo et al., 2023; Liu et al., 2024a).
@@
$\bullet$ Model architecture. For the model architecture, Flan-T5 (Chung et al., 2024) based on the encoder-decoder architecture has almost no ability to recommend items in the cold-start setting, as its training corpus does not involve specialized instructions of our task. Trained with prompts of recommendations, the encoder-decoder architecture is suitable for prompt tuning and instruction tuning (Zhang et al., 2023e; Geng et al., 2022; Chu et al., 2023). Similarly, the first and first version of ChatGLM (Zeng et al., 2023) performs poorly on the zero-shot ranking task, and is not as good as Vicuna and LLaMA2 based on the causal decoder. However, the third version of ChatGLM, i.e., ChatGLM3 has comparable recommendation performance with Vicuna (Chiang et al., 2023) and LLaMA2 (Touvron et al., 2023b), which further indicates the importance of selecting an advanced foundation model. In terms of the series of LLaMA models (Touvron et al., 2023a), LLaMA3 achieves the best performance, and Vicuna is better than LLaMA, which is related to their training data and release time. Furthermore, the chat version of LLaMA2, i.e., LLaMA2-chat is a series that uses conversational dialog instructions to fine-tune LLaMA2 (Touvron et al., 2023b), which is more suitable for our tasks in ranking settings. Therefore, the results of LLaMA2-chat are significantly improved compared with LLaMA2, and LLaMA2-70B-chat and LLaMA3-70B achieve better performance than the closed-source model ChatGPT. Generally speaking, researchers prefer to study recommendation tasks based on the causal decoder framework such as LLaMA, and the latest version of LLaMA has a better generalization ability.
@@
(b) Amazon Books
Figure 4. The recommendation performance and inference time of LLMs w.r.t. the parameter scale. We compare the different parameter scales of LLMs, and the larger the symbol in the figure, the larger the parameter scale.
[…]
The recommendation performance and inference time of LLMs w.r.t. the parameter scale. We compare the different parameter scales of LLMs, and the larger the symbol in the figure, the larger the parameter scale.